MULTISTAGE VIBRATION - TORSION - ROTATION COUPLING IN THE ν_{6} ASYMETRIC N-O STRETCH OF NITROMETHANE

\underline{LOU DEGLIUMBERTO}, DAVID S. PERRY, University of Akron, Akron, OH 44325; WENDY HEISERMAN, REBECCA COOK, HOWARD D. METTEE , Youngstown State University, Youngstown, OH 44555; ROBERT L. SAMS, Pacific Northwest National Laboratory, Richland, WA 99352; GEORG O. SORENSEN^{a}, University of Copenhagen, DK-2100, Copenhagen, Denmark.

In the high resolution spectrum of nitromethane near 1584 cm^{−1}each of the zeroth order transitions of the lowest internal rotor state is split into a clump of 4 or 5 rovibrational transitions. Individual perturbing (dark) states are followed as a function of the J ′, K′_{a}, and K′_{c} rotational quantum numbers. Four stages of vibration-torsion-rotation coupling are identified: (i) sub-picosecond c-axis Coriolis coupling to a torsionally excited state dark vibration (ν_{7} + ν_{10} or ν_{5}), (ii) a-axis Coriolis coupling between A′_{1} and A′_{2} torsional states of the dark vibration, (iii) c-axis Coriolis coupling of the dark vibration back to dark rotational states of the asymmetric N-O stretch, and (iv) at J ′=8 and higher a strong coupling to a second torsionally excited dark vibration .

adeceased